OpenStudy (anonymous):

X^2 + 2x - 3 = x(x + 2) - 3 Is this a factoring error?

hero (hero):

There's a specific procedure for factoring quadratic expressions. The above is not a correct approach.

hero (hero):

For a quadratic expression in the form ax^2 + bx + c, if it is factorable, the general method is as follows: `1.` Find two numbers that multiply to get ac, yet add to get b. `2.` Split the middle term bx, using those two numbers `3.` Factor the first two terms and the last two terms by grouping `4.` Factor out the remaining common term

OpenStudy (anonymous):

Thanks Hero! Why does the approach in my question not work?

hero (hero):

For x^2 + 2x - 3: `1.` The two numbers are 3 and -1, since ac = (3)(-1) = -3 and b = 3 - 1 = 2 `2.` Splitting the middle term you get: x^2 + 3x - 1x - 3 `3.` Factoring by grouping you get: (x^2 + 3x) + (-1x - 3) x(x + 3) - 1(x + 3) `4.` x + 3 is the remaining common term, so that is factored out to get: (x + 3)(x - 1)

hero (hero):

So the algebraic steps alone look like this: x^2 + 2x - 3 Original Problem = x^2 + 3x - 1x - 3 Subtraction Property (2x = 3x - 1x) = x(x + 3) - 1(x + 3) Factor by Grouping = (x + 3)(x - 1) Distributive Property ab + ac = a(b + c)

hero (hero):

Some students are able to factor in their heads and do it all in one step, going from the original quadratic to the final result: x^2 + 2x - 3 = (x + 3)(x - 1) It's a very intuitive method that requires use of mental math and mastery of multiplication and addition/subtraction.